OpenStudy (jpsmarinho):

How to solve this: integrate 1/(x^2+9)dx?

OpenStudy (anonymous):

look in the back cover of your calc book for integrals of the form \[\int\frac{dx}{a^2+x^2}\] and you will see that it is \[\frac{1}{a}\tan^{-1}(\frac{x}{a})\]

OpenStudy (anonymous):

you know \[\Large \frac{d}{dx}(\tan^{-1}(x)=\frac{1}{x^2+1}\]

OpenStudy (anonymous):

guys mind if I watch this?

OpenStudy (anonymous):

or recall that \[\frac{d}{dx}[\tan^{-1}(x)]=\frac{1}{1+x^2}\] and adjust for the constant

OpenStudy (anonymous):

what @sami-21 said

OpenStudy (jpsmarinho):

Thanks!